OpenStudy (anonymous):

Find the exact value of sin-1the quantity square root of three divided by two?

OpenStudy (ranga):

\[\Large sin^{-1}\frac{\sqrt(3)}{2} = ?\]Look up the unit circle or the trig table to see for what angle theta is sin(theta) = sqrt(3)/2.

OpenStudy (anonymous):

60 degrees?

OpenStudy (ranga):

Yes. 60 degrees is your answer. If they want the answer in radians it is pi/3 radians.

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (ranga):

You are welcome.